OpenStudy (anonymous):

Write y = –0.5x + 0.8 in standard form using integers.

OpenStudy (anonymous):

Turn -0.5 and 0.8 into a fraction, which would be \[ y = -\frac{1}{2} + \frac{4}{5}\] Now find a common denominator for the two fraction and times it on both sides \[ 10*y = 10 * [-\frac{1}{2} + \frac{4}{5}]\] Once you do the calculation, put it into standard form. Can you finish it?